How much cacao must be added to 12kgs of chocolate containing 50% cacao in order that the resulting mixture will contain 75% cacao?
Answer by Earlsdon(6103)

Let x = the number of kgs of cacao to be added to the 12 kgs of 50% cacao to obtain (12+x)kgs of a mixture containing 75% cacao.
This can be expressed by:
x+12(0.5) = (12+x)(0.75) Simplify and solve for x.
x+6 = 9+0.75x Subtract 0.75x from both sides.
0.25x+6 = 9 Subtract 9 from both sides.
0.25x = 3 Finally divide both sides by 0.25
x = 12
12 kgs of cacao must be added.